A bag contains 200items of which 20items are defectives.seven items are selected at random 1 aftert another with replacement.what is the probability of obtaning ...questions number 1)no defective 2)5non defective

from: Probability distribution of binomial problem.

Answer by stanbon(75887)

A bag contains 200 items of which 20 items are defectives. seven items are selected at random 1 after another with replacement. what is the probability of obtaning ...questions number
1)no defective
Binomial Problem with n = 7 and p(defect) = 20/200 = 1/10
---
P(x = 0) = 7C0(0.1)^0*(0.9)^7 = 0.9^7 = 0.4783
-----------------------------------------------------
2)5 non defective
that is the same as 2 defective out of 7
P(x = 2) = 7C2(0.1)^2*(0.9)^5 = 0.1240
=================
Cheers,
Stan H.
=================
--------------------